Circular No. 7680

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001dh, (2001dj) and 2001dl F. Patat, European Southern Observatory; C. Contreras and J. Prieto,

Universidad Catolica de Chile; and G. Altavilla, S. Benetti, E. Cappellaro, A. Pastorello and M. Turatto, Osservatorio Astronomico di Padova, report: “Spectra of three candidate supernovae were obtained on Aug. 8 using the Danish 1.5-m telescope + DFOSC (spectral range 320–900 nm, resolution 1.2 nm) at La Silla. 2001dh (IAUC 7670) is a type-II supernova, ∼ 1 month after maximum, the spectrum being dominated by the Hα P-Cyg line with the absorption minimum measured at 641 nm and yielding an expansion velocity of ∼ 9000 km/s, assuming the NED recession value of 2608 km/s for the parent galaxy; other strong lines are Hβ (478 nm), Fe II (513 and 497 nm) and Ca II (842 nm). 2001dj (IAUC 7674) is not a supernova but most likely a galactic variable of early-F spectral type. 2001dl (IAUC 7675) is a type-Ia supernova, its spectrum being very similar to that of SN 1994D at maximum (cf. Patat et al. 1996, Monthly Notices Roy. Astron. Soc. 278, 111); the characteristic Si II absorption is measured at 627 nm, and given the NED recession velocity of 6204 km/s for the parent galaxy, the expansion velocity is as large as 10 000 km/s.”

WZ SAGITTAE T. Kato, H. Ohashi, R. Ishioka, M. Uemura and K. Matsumoto, Kyoto

University; G. Masi and F. Mallia, Ceccano, Italy; D. Starkey, Auburn, Indiana; J. Pietz, Erftstadt, Germany; B. Martin, Edmonton, Alberta; G. A. Good, Albuquerque, New Mexico; and M. Richmond, S. Davis and T. Davis, Rochester, New York, on behalf of the VSNET collaboration team, report: “Our latest time-resolved CCD photometry (IAUC 7678) confirmed that there was an interchange of the main maxima from early to ‘genuine’ superhumps around Aug. 5.8 UT. The signal of superhumps further evolved, and their times of maxima during Aug. 5.25–11.12 are well represented by HJD(max) = 2452126.755 + 0.057153(16)E. The residuals from this ephemeris show a systematic variation with probable period 4 days. This variation is most likely due to the beat between the orbital and superhump periods. On Aug. 9 the superhumps showed prominent ‘secondary superhumps’ at phase 0.4, and these became even stronger on Aug. 10, gradually merging into singly-peaked superhumps on Aug. 11.1. The present finding concerning the evolution and period of superhumps provides higher-quality confirmation of the interpretation of the early behavior of the 1978 outburst proposed by Bohusz and Udalski (IBVS 1583).” Masi provides the following improved position for WZ Sge (10 USNO-A2.0 reference stars): α = 20^{h}07^{m}36^{s}.53, δ = +17^{o}42′15′′.3 (J2000.0, mean residual 0′′.25).

2001 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden